give the dimensions of five rectangles that have a perimeter o 50 feet
Answer by checkley77(12844)

P=2L+2W
50=2*10+2*15=20+30=50
50=2*5+2*2*20=10+40=50
50=2*4+2*21=8+42=50
50=2*24+2*1=48=2=50
50=2*3+2*22=6+44=50